A new surgery is successful 85% of the time. If the results of 7 such surgeries are randomly sampled, what is the probability that at least 6 of them are successful? Carry your intermediate computations to at least four decimal places, and round your answer to two decimal places.
Answer #1

X ~ Bin ( n , p )

Where n = 7 , p = 0.85

Binomial probability distribution is

P(X) = nCx px ( 1- p)n-x

So,

P(X >= 6) = P(X = 6) + P(X = 7)

= 7C6 * 0.856 * (1 - 0.85) + 7C7 * 0.857 * (1 - 0.85)0

= 0.72